Eiwoh2:

Which of the following is a polynomial with roots: −square root of 3, square root of 3 , and 2? x3 + 3x2 − 5x − 15 x3 + 2x2 − 3x − 6 x3 − 3x2 − 5x + 15 x3 − 2x2 − 3x + 6

Eiwoh2:

@dude @Shadow I require A S S I S T A N C E ;-;

AngeI:

Not my type of math sorry

Eiwoh2:

Lol it's all right. I appreciate the effort of coming onto the question lolol XD

dude:

To find the roots of an equation set it equal to 0

dude:

(I assume you can use a calculator but yeah, uh it’s not A)

Vocaloid:

factor out each equation let's use B as an example x3 + 2x2 − 3x − 6 you can factor the first two terms into (x^2)(x+2) and the second two terms into -3(x+2) giving us (x^2)(x+2) -3(x+2) notice that you have an (x+2) part in each of these two terms, so you can also factor out (x+2) to get (x^2-3)(x+2) = 0 setting each component equal to 0 gives us (x^2-3) = 0 , x = +/- sqrt(3) (x+2) = 0, x = -2 so B isn't quite it but you can repeat the process with the other choices to see what the right solution is